OpenStudy (anonymous):

17. A tree casts a shadow of 20ft. The angle of elevation from the tip of the shadow to the top of the tree is 70 degrees. To the nearest foot, what is the height of the tree?

OpenStudy (anonymous):



OpenStudy (anonymous):

Can you solve for \(x\)?

OpenStudy (anonymous):

Would it be 25? @wio

OpenStudy (anonymous):

@HeyImKae Tell me how you did it, and I'll tell you if it is corect.

OpenStudy (anonymous):

I entered 70 * tan20 into my calculator! @wio

OpenStudy (anonymous):

The angle and side are mixed up...

OpenStudy (anonymous):

so would it be 20 * tan70? Because then I get 54

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

that sounds about right.

OpenStudy (anonymous):

54 ft tree.

jhonyy9 (jhonyy9):

so tan70 = x/20 than x = 20 *tan70 = ?